a) Two parallel plates with 0.025m width and 0.05m length are placed vertically such that the field intensity E between them is 11kN/C directed upward. Find the charge of each plate.

b) A metal sphere with radius 0.035m has a charge 5 x10-6C at its surface. Find the electric field E at the surface, the field at 0.02m out of the surface and the field at 0.02m within the surface.



1
Expert's answer
2021-10-07T08:49:17-0400

a)


q=ϵEAq=(8.851012)(11000)(0.050.025)=1.21010 Cq=\epsilon EA\\q=(8.85\cdot10^{-12})(11000)(0.05\cdot0.025)\\=1.2\cdot10^{-10}\ C

b)


E0=(9109)(5106)(0.035)2=3.7107VmE1=(9109)(5106)(0.035+0.02)2=1.5107VmE2=0VmE_0=(9\cdot10^{9})\frac{(5\cdot10^{-6})}{(0.035)^2}=3.7\cdot10^{7}\frac{V}{m}\\ E_1=(9\cdot10^{9})\frac{(5\cdot10^{-6})}{(0.035+0.02)^2}=1.5\cdot10^{7}\frac{V}{m}\\ E_2=0\frac{V}{m}